# Insights into the structure and kinematics of a Milky Way-like galaxy

### v2

@@ 3.2 Stellar Kinematics
Figure 3: Top-view representations of the tangential ($V_{\phi}$, left panel) and radial ($V_{R}$, right panel) velocity fields for the stellar component of our model. Note that $V_{\phi}$ is negative due to the clockwise rotation of the galaxy. The Sun is located at a viewing angle of $\phi_{\mathrm{obs}}=20^{\circ}$ with respect to the bar. The original Sun’s position as per the best viewing angle from DuranCamacho2024 is marked as a black circle, and the Sun’s reflection point (SunRP) at $180^{\circ}$ offset is marked with a cross.
@@
It is well known that the motions of stars and gas in galactic bars follow a quadrupole pattern, and this pattern has also been observed in the MW’s Galactic Centre (e.g. with Gaia DR2 and EDR3, Bovy2019; queiroz2021milky). One of the sharpest views of this kinematical pattern was revealed by drimmel2023, who analysed the kinematics of RGB and OB stars in the Milky Way using data from Gaia DR3. They found strong non-circular motions associated with this quadrupole pattern, with mean radial motions of the order of $\pm 40$ km s-1. They also showed that streaming motions occur at larger Galactocentric radii ($R>5$ kpc), with possible association with Lindblad resonances. For the tangential component, they observed a clear elongated feature with lower tangential velocities along the bar’s major axis, within a $5$ kpc radius.
@@
In order to benchmark our model against those observations, and understand how closely our model is able to replicate the kinematics of the centre of the MW, here we investigate the velocity fields associated with the bar pattern. In Fig. 3 we show the tangential ($V_{\phi}$, left) and radial ($V_{R}$, right) velocities of our model’s stellar disc, up to a radius of $12$ kpc. In this Figure, the bar is aligned with the $y=0$ axis, and we position the Sun at a distance of $8.275$ kpc from the galactic centre (GRAVITYCollaboration2021), and such that the galactic bar has an inclination angle of $\sim 20^{\circ}$ with respect to the observer, to match the results from drimmel2023333This angle was also found to be a suitable position for the observer in our original work (DuranCamacho2024).
@@
The radial velocity component in the right panel of Fig. 3 exhibits a quadrupole pattern within the long-bar extent, i.e. $\sim 5$ kpc from the galactic centre. The radial velocities reach values of $\pm 45$ km s-1, which is remarkably consistent with observed findings. As shown on the left panel of Fig. 3, the tangential velocity component also reveals an extended feature along the bar’s major axis with lower velocities, as expected, given that the stars are in $x_{1}$ orbits, with circular velocities closer to the bar pattern speed. For our model, those velocities are of the order of $\sim 40-70$ km s-1 within the inner 5 kpc, which is similar to the observations.
@@
As mentioned in DuranCamacho2024, our simulation does not develop the $x_{2}$ orbits typically observed in the vicinity of the Central Molecular Zone, which has an extension of $\sim 250$ pc around the Galactic Centre (see e.g. Henshaw2016). These orbits are oriented perpendicular to the Galactic bar and are expected within the ILR for non-axisymmetric potentials (e.g. Contopoulus1989; athan92a). They emerge from the interplay between the bar’s dynamics and the stellar and mass profile of the bulge. These type of orbits may not develop if there is a lack of ILRs, a lack of mass in the inner regions (e.g. Hasan1993; Athanassoula2013), or the presence of a very strong bar potential (e.g. Contopoulus1989; athan92a). The existence of $x_{2}$ orbits is also sensitive to the pattern speed of the bar, which determines the position of the ILRs. Faster bars push the ILR inwards, while slower bars place the ILR farther out, allowing for a more extended $x_{2}$ region (see e.g. athan92a; sellwood1993). Given that our model’s pattern speed ($30\pm 0.2$ km s-1 kpc-1) is comparable to the observed values (304030-40 km s-1 kpc-1, see e.g. drimmel2023; clarke2022), and that within our suite of models from DuranCamacho2024, some of the simulations that do develop $x_{2}$ orbits have very similar pattern speeds, we conclude that the pattern speed is not the defining factor in this case. In addition, in DuranCamacho2024, we demonstrated that our model does feature ILRs at $\sim$0.2 and $\sim$1.1 kpc, and thus that is not the reason the orbits do not develop.
@@
Figure 4: Spherical enclosed mass as a function of Galocentric radius for the analytical model from sormani2022stellar (blue-dashed line), and our model at initial (orange-dotted line), optimal (green-dotted line), and late (red-dotted line) times. The position of the two ILRs in our model are shown as green vertical dashed lines.
@@
A sufficiently high central mass concentration is known to contribute to the support of $x_{2}$ orbits (athan92a; binney1991understanding; Hasan1993). In order to understand if the lack of $x_{2}$ orbits in our model could indeed be a result of mass missing in the inner parts of the model (as potentially hinted by Sect. 3.1), here we investigate the total (spherically) enclosed stellar mass as a function of galactocentric radius $R$, for our model at different times, as well as the S22444Note that although here we show just the comparison to S22, the results would be near identical if using the R17 model instead, given that the two profiles are a close match to each other, for $R_{\rm gal}<2.5$ kpc (in $xy$), and $|z|<1$ kpc (see Fig.12). analytical model (Fig.4)
@@
In the previous section we found the largest discrepancies in the surface density distributions in the $z$ direction, and hence we also investigate the enclosed mass as a function of vertical direction $z$ (Appendix A). We find that in the mid-plane (within $|z|$ ¡ 250 pc), our model is not missing any mass until well beyond the second IRL, and that indeed it is only at higher $z$ distances that the differences become noticeable. Given that $x_{2}$-type of orbits develop in the mid-plane (e.g. athan92a; pichardo2002), it is unlikely that the lack of stellar mass in the inner galaxy of our models is the main contributing factor to the absence of these orbits. This is also corroborated by the fact that despite the similarities in the S22 and R17 stellar profiles for $R<2.5$ kpc, the hydrodynamical simulations from Portail2015 from which the S22 profile is based, also do not develop the $x_{2}$ orbits, while those from R17 do (see also sormani2019geometry; tress2020-CMZ).
An alternative explanation could thus be linked to the strength of the galactic bar itself. athan92a demonstrated that the prevalence of these orbits diminishes as the strength of the bar increases. Consequently, it could be that our model’s bar, by being effectively “narrower” than the observed one, might mean that the relative bar strength in our simulation is sufficiently high to suppress the formation of $x_{2}$-type orbits. It is possible that the inclusion of stellar and supernova feedback into these models will help steer up the dynamics of the gas and stars, and help diminish the relative strength of the bar potential - thus potentially facilitating the development of these $x_{2}$-type orbits. This will be investigated in follow up work.
@@ 4.1 Spiral pattern
We use Fourier transforms for this purpose, as previously applied in nearby galaxies observations (e.g. Elmegreen1989; Rix1995; Fuchs1999) and numerical simulations (e.g. Bottema2003). By decomposing the observed structures of the galaxy into their harmonic components, this method serves as an indicator of the multiplicity/periodicity of structures in a galactic disc. The interpretation of the results from this method often relies on the assumption that the galaxy’s spiral structure can be approximated by a periodic function in the azimuthal direction, with the Fourier coefficients providing a direct link to the underlying symmetry and number of arms. More details on this technique can be found in Appendix B.
The final representation of our model’s Fourier amplitudes, $C_{n}$, of each Fourier coefficient (with harmonic number $n$), is shown in Fig. 5 as density plots across radial bins. The amplitude of each coefficient indicates the prominence of that type of multiplicity, and therefore, a prominent peak mode $n$ could be seen as directly linked to the dominant number $n$ of spiral arms at a given radius. However, as pointed out by Elmegreen1989, caution should be taken when interpreting the dominant mode as indicative of the exact number of spiral arms, as there is some degeneracy. As such, we use this Fourier decomposition as a means to inter-compare the multiplicity of the stellar and gaseous distributions, rather than to quantify their exact number of arms.
@@
For the stellar component (Fig. 5, top panel), we find a dominant $n=2$ configuration extending up to $\sim 7$ kpc, with the highest amplitudes near the bar. Beyond $\sim 8$ kpc, the pattern weakens and transitions to multiple components as the stellar potential becomes more diffuse, with no particular dominant mode. This is also illustrated in Figure 6, where we show the dominant mode $n$ as a function of radius, and we can see that the stars maintain a dominant $n=2$ mode up to $\sim 7$ kpc before becoming more chaotic. These findings align with observations in external galaxies, where inner discs often show 2-arm structures, while outer regions display a more flocculent nature. We also detect a much weaker $n=4$ mode, which could be indicative of a 2+2 arm pattern, where two strong arms are accompanied by two weaker arms, consistent with Gaia studies suggesting a similar Milky Way structure (Poggio2021).
@@
These findings are consistent with the complexity of the spiral patterns that we find when extracting the spiral arm ridges from the sharpened maps of the top-down stellar and gas surface densities (see Fig. 15). The stellar spiral arms, are typically low contrast and smoother in nature (hence not being particularly visually prominent in Figure 1), and we find $\sim 2-4$ arm segments in any radial cross section from 3 kpc up to 12 kpc. In the gas, however, the spiral pattern is sharper, which is consistent with observations of external galaxies, where spiral structure is often more apparent in the gas, dust, and young stars than in the old stellar disc (e.g. Elmegreen1987; Elmegreen2011). That said, our model’s spiral arm ridges in the gas are significantly more numerous and segmented than the underlying stellar pattern. On average, we find $\sim 5-8$ gaseous ridges in any radial cross section from 3 kpc up to 12 kpc. In other words, the gas spiral structure shows significant segmentation into smaller arm sections rather than following a well-defined large-scale coherent spiral pattern (see also Fig. 9).
Overall, our model does not conform to a grand design spiral archetype. Instead, while a spiral pattern is always present, it is mostly transient, with the gas and stars often out of sync. This mirrors observations of external galaxies, where gaseous spirals display greater variability due to the interstellar medium’s dynamical response (see e.g. Elmegreen1987; Maschmann2024; Dobbs2008; Renaud2013). Such hybrid behaviour highlights the diversity of spiral structures under evolving galactic conditions (Foyle2010; Hart2017).
As noted in Sect. 1.2, studies suggest the Milky Way shows both grand design and multi-armed features. For example, drimmel2001 and churchwell2009 found a two-arm structure in the inner Galaxy, while the outer disc is more complex. Spitzer GLIMPSE observations confirm this duality, showing dominant and fainter arms (benjamin2005). Gas surveys (Nakanishi2003; kalberla2009hi) also reveal significant variation, consistent with our model. This suggests that the Milky Way’s spiral structure may be inherently transient, with gas responding in a complex, non-density-wave-like manner, a point we explore further in the following sections (4.2 and 5.1).
@@ 4.2 Solar neighbourhood
The recent work by khanna2023measuring has identified significant streaming motions for stars within $3.5$ kpc from the Sun, demonstrating tentative correlations between the velocity structures and the spiral arms. Using an axisymmetric model to subtract from the observed velocities, they quantified overall velocity trends and found that gradients in the tangential and radial components ($\Delta(-V_{\phi}$) and $\Delta V_{R}$) were not uniform, with variations averaging $\sim 5$ km s-1 and $\sim 10$ km s-1 respectively. However, the extent to which such velocity patterns can be related to the dynamics of the stellar and gaseous spiral pattern in the MW is incredibly hard to ascertain from observations. Indeed, as mentioned in Section 1.2, our knowledge of the position and extent of the stellar and gaseous arms remains highly uncertain, making these results tentative at best.
@@
In order to shed some light into how the observed stellar kinematics pattern from Gaia may inform us of the underlying spiral arm structure and kinematics in the solar neighbourhood, here we investigate the equivalent stellar velocity patterns in our model, and then compare that to the underlying stellar and gaseous spiral pattern. For that, we obtain the velocity residuals in a similar way to the observations by khanna2023measuring by contrasting the actual velocity fields observed in our simulation against an idealised rotation curve (see Appendix D for more details).
@@
In Figure 7, we present the heliocentric velocity residuals, $\Delta(-V_{\phi})$ and $\Delta V_{R}$, for the data from the Gaia DR3 study by khanna2023measuring (top), alongside our model’s stellar (middle) and gaseous (bottom) components within the mid-plane (—zz— ¡ 0.25 kpc). On the Gaia data maps we include the spiral arm tracks from reid19, and in the simulation panels, we have included the stellar and gaseous spiral arm ridges (in black and grey respectively), as identified using an unsharpening technique (see full details in Appendix C). From this Figure, we can see that the model’s gaseous spiral ridges loosely follow the tracks suggested by reid19, and that the general trends for the stellar kinematics in our model seem to be similar to those observed by khanna2023measuring, with the sun located at the interface between inward and outward motions, and in a region of moderately slower tangential velocities. Both our model and the observations show strong gradients in $\Delta(-V_{\phi})$ and $\Delta\mathrm{V}_{\mathrm{R}}$, albeit slightly more pronounced in the model, with magnitudes reaching as high as $15$ km s-1 and $25$ km s-1 for the tangential and radial components respectively. Overall, the stellar kinematics in our simulation show a relatively weak correlation with the positions of the spiral ridges (be those the gaseous or stellar arms), as none of the velocity components demonstrate a clear systematic alignment.
@@
Figure 7: Heliocentric velocity residual maps for $\Delta(-V_{\phi})$ (left) and $\Delta V_{R}$ (right) for Gaia DR3 data from khanna2023measuring (top), and our model’s stellar (middle) and gaseous (bottom) components in the galactic mid-plane (—zz—¡ 0.25 kpc). The stellar maps use a binning identical to those of the Gaia data, of $250\times 250$ pc2 pixels, for a direct comparison, while the gas maps have been binned to cells of $60\times 60$ pc2. The solid grey lines on the top panels represent the gaseous spiral arm tracks from reid19. For our simulation, the stellar and gaseous spiral arm ridges extracted from the sharpened images of Fig. 15, are superimposed in black and grey respectively. The position of the sun is marked as a black circle, and the galactic centre is outside this image, at (8.275,0) kpc.
@@
Overall, our analysis shows that the pattern in the stellar structure and kinematics of our model at the chosen time is generally mild and consistent with Gaia observations of RGB stars, but that it translates into a gaseous pattern that is much sharper. This is also consistent with observations from drimmel2023 who show that OB stars (which by being younger, are potentially better tracers of the underlying gas distribution that originated them) do align better with the gaseous spiral structures, specially towards the inner disc, but with low number densities their maps remain uncertain. Given the consistency between our results and MW observations, our model could therefore offer some insight into how the formation (and evolution) of the gaseous spiral structures of the MW, albeit potentially rapidly evolving, might be intricately inter-connected with the observed kinematics of the gas, and thus we explore this correlation further in the next section.

### v3

@@ 3.2 Stellar kinematics
Figure 3: Top-view representations of the tangential ($V_{\phi}$, left panel) and radial ($V_{R}$, right panel) velocity fields for the stellar component of our model. Note that $V_{\phi}$ is negative due to the clockwise rotation of the galaxy. The bar is aligned with the $y=0$ axis and the Sun is located at a viewing angle of $\phi_{\mathrm{obs}}=20^{\circ}$ with respect to the bar. The original Sun’s position as per the best viewing angle from Durán-Camacho et al. (2024) is marked as a black circle, and the Sun’s reflection point (SunRP) at $180^{\circ}$ offset is marked with a cross.
@@
It is well known that the motions of stars and gas in galactic bars follow a quadrupole pattern, and this pattern has also been observed in the MW’s Galactic Centre (e.g. with Gaia DR2 and EDR3, Bovy et al., 2019; Queiroz et al., 2021). One of the sharpest views of this kinematical pattern was revealed by Gaia Collaboration et al. (2023), who analysed the kinematics of RGB and OB stars in the Milky Way using data from Gaia DR3. They found strong non-circular motions associated with this quadrupole pattern, with mean radial motions of the order of $\pm 40$ km s-1. They also showed that streaming motions occur at larger Galactocentric radii ($R>5$ kpc), with possible association with Lindblad resonances. For the tangential component, they observed a clear elongated feature with lower tangential velocities along the bar’s major axis, within a $5$ kpc radius.
@@
In order to benchmark our model against those observations, and understand how closely our model is able to replicate the kinematics of the centre of the MW, here we investigate the velocity fields associated with the bar pattern. In Fig. 3 we show the tangential ($V_{\phi}$, left) and radial ($V_{R}$, right) velocities of our model’s stellar disc, up to a radius of $12$ kpc. In this Figure, the bar is aligned with the $y=0$ axis, and we position the Sun at a distance of $8.275$ kpc from the galactic centre (GRAVITY Collaboration et al., 2021), and such that the galactic bar has an inclination angle of $\sim 20^{\circ}$ with respect to the observer, to match the results from Gaia Collaboration et al. (2023) [^3] This angle was also found to be a suitable position for the observer in our original work (Durán-Camacho et al., 2024).
@@
The radial velocity component in the right panel of Fig. 3 exhibits a quadrupole pattern within the long-bar extent, i.e. $\sim 5$ kpc from the galactic centre. This quadrupole arises from bar-driven non-circular streaming motions: the sign of $V_{R}$ alternates across the bar major and minor axes (yy=0 and xx=0 lines respectively in Fig. 3) as stars follow elongated bar-supporting orbits, producing approaching/receding (inward/outward) flows within the bar region. The radial velocities reach values of $\pm 45$ km s-1, which is remarkably consistent with observed findings. As shown on the left panel of Fig. 3, the tangential velocity component also reveals an extended feature along the bar’s major axis with lower velocities, as expected, given that the stars are in $x_{1}$ orbits, with circular velocities closer to the bar pattern speed. For our model, those velocities are of the order of $\sim 40-70$ km s-1 within the inner 5 kpc, which is similar to the observations.
@@
As mentioned in Durán-Camacho et al. (2024), our simulation does not develop the $x_{2}$ orbits typically observed in the vicinity of the Central Molecular Zone, which has an extension of $\sim 250$ pc around the Galactic Centre (see e.g. Henshaw et al., 2016). These orbits are oriented perpendicular to the Galactic bar and are expected within the ILR for non-axisymmetric potentials (e.g. Contopoulos and Grosbol, 1989; Athanassoula, 1992). They emerge from the interplay between the bar’s dynamics and the stellar and mass profile of the bulge. These type of orbits may not develop if there is a lack of ILRs, a lack of mass in the inner regions (e.g. Hasan et al., 1993; Athanassoula et al., 2013), or the presence of a very strong bar potential (e.g. Contopoulos and Grosbol, 1989; Athanassoula, 1992). The existence of $x_{2}$ orbits is also sensitive to the pattern speed of the bar, which determines the position of the ILRs. Faster bars push the ILR inwards, while slower bars place the ILR farther out, allowing for a more extended $x_{2}$ region (see e.g. Athanassoula, 1992; Sellwood and Wilkinson, 1993). Given that our model’s pattern speed ($30\pm 0.2$ km s-1 kpc-1) is comparable to the observed values (304030-40 km s-1 kpc-1, see e.g. Gaia Collaboration et al., 2023; Clarke and Gerhard, 2022), and that within our suite of models from Durán-Camacho et al. (2024), some of the simulations that do develop $x_{2}$ orbits have very similar pattern speeds, we conclude that the pattern speed is not the defining factor in this case. In addition, in Durán-Camacho et al. (2024), we demonstrated that our model does feature ILRs at $\sim$0.2 and $\sim$1.1 kpc, and thus that is not the reason the orbits do not develop.
@@
Figure 4: Spherical enclosed mass as a function of Galocentric radius for the analytical model from Sormani et al. (2022) (blue-dashed line), and our model at initial (orange-dotted line), optimal (green-dotted line), and late (red-dotted line) times. The position of the two ILRs in our model are shown as green vertical dashed lines.
@@
A sufficiently high central mass concentration is known to contribute to the support of $x_{2}$ orbits (Athanassoula, 1992; Binney et al., 1991; Hasan et al., 1993). In order to understand if the lack of $x_{2}$ orbits in our model could indeed be a result of mass missing in the inner parts of the model (as potentially hinted by Sect. 3.1), here we investigate the total (spherically) enclosed stellar mass as a function of galactocentric radius $R$, for our model at different times, as well as the S22444Note that although here we show just the comparison to S22, the results would be near identical if using the R17 model instead, given that the two profiles are a close match to each other, for $R_{\rm gal}<2.5$ kpc (in $xy$), and $|z|<1$ kpc (see Fig.12). analytical model (Fig.4)
@@
In the previous section we found the largest discrepancies in the surface density distributions in the $z$ direction, and hence we also investigate the enclosed mass as a function of vertical direction $z$ (Appendix A). We find that in the mid-plane (within $|z|$ ¡ 250 pc), our model is not missing any mass until well beyond the second IRL, and that indeed it is only at higher $z$ distances that the differences become noticeable. Given that $x_{2}$-type of orbits develop in the mid-plane (e.g. Athanassoula, 1992; Pichardo et al., 2002), it is unlikely that the lack of stellar mass in the inner galaxy of our models is the main contributing factor to the absence of these orbits. This is also corroborated by the fact that despite the similarities in the S22 and R17 stellar profiles for $R<2.5$ kpc, the hydrodynamical simulations from Portail et al. (2015) from which the S22 profile is based, also do not develop the $x_{2}$ orbits, while those from R17 do (see also Sormani et al., 2019; Tress et al., 2020).
An alternative explanation could thus be linked to the strength of the galactic bar itself. Athanassoula (1992) demonstrated that the prevalence of these orbits diminishes as the strength of the bar increases. Consequently, it could be that our model’s bar, by being effectively “narrower” than the observed one, might mean that the relative bar strength in our simulation is sufficiently high to suppress the formation of $x_{2}$-type orbits. It is possible that the inclusion of stellar and supernova feedback into these models will help steer up the dynamics of the gas and stars, and help diminish the relative strength of the bar potential - thus potentially facilitating the development of these $x_{2}$-type orbits. This will be investigated in follow up work.
@@ 4.1 Spiral pattern
We use Fourier transforms for this purpose, as previously applied in nearby galaxies observations (e.g. Elmegreen et al., 1989; Rix and Zaritsky, 1995; Fuchs and Möllenhoff, 1999) and numerical simulations (e.g. Bottema, 2003). By decomposing the observed structures of the galaxy into their harmonic components, this method serves as an indicator of the multiplicity/periodicity of structures in a galactic disc. The interpretation of the results from this method often relies on the assumption that the galaxy’s spiral structure can be approximated by a periodic function in the azimuthal direction, with the Fourier coefficients providing a direct link to the underlying symmetry and number of arms. More details on this technique can be found in Appendix B.
The final representation of our model’s Fourier amplitudes, $C_{n}$, of each Fourier coefficient (with harmonic number $n$), is shown in Fig. 5 as density plots across radial bins. The amplitude of each coefficient indicates the prominence of that type of multiplicity, and therefore, a prominent peak mode $n$ could be seen as directly linked to the dominant number $n$ of spiral arms at a given radius. However, as pointed out by Elmegreen et al. (1989), caution should be taken when interpreting the dominant mode as indicative of the exact number of spiral arms, as there is some degeneracy. As such, we use this Fourier decomposition as a means to inter-compare the multiplicity of the stellar and gaseous distributions, rather than to quantify their exact number of arms.
@@
For the stellar component (Fig. 5, top panel), we find a dominant $n=2$ configuration extending up to $\sim 4$ kpc, with the highest amplitudes near the bar. Beyond that point, the pattern weakens and transitions to multiple components as the stellar potential becomes more diffuse, with no particular dominant mode. This is also illustrated in Figure 6, where we show the dominant mode $n$ as a function of radius, and we can see that the stars maintain a dominant $n=2$ mode up to $\sim 4$ kpc before becoming more chaotic. These findings align with observations in external galaxies, where inner discs often show 2-arm structures, while outer regions display a more flocculent nature. We also detect a much weaker $n=4$ mode, which could be indicative of a 2+2 arm pattern, where two strong arms are accompanied by two weaker arms, consistent with Gaia studies suggesting a similar Milky Way structure (Poggio et al., 2021).
@@
These findings are consistent with the spiral patterns inferred from the spiral-ridge extraction applied to the sharpened top-down stellar and gas surface density maps (Fig. 15). In the stellar component, the ridges are low-contrast and comparatively smooth (and therefore not particularly prominent by eye in Fig. 1); nevertheless, in any radial cross section we identify $\sim 2$–4 stellar arm segments over $R\simeq 3$–12 kpc. In the gas, however, the spiral pattern is sharper, in line with observations of external galaxies, where spiral structure is often more apparent in the gas/dust, and young stars than in the old stellar disc (e.g. Elmegreen and Elmegreen, 1987; Elmegreen et al., 2011). That said, the gaseous spiral pattern is also more fragmented: we typically identify $\sim 5$–8 gaseous ridges per radial cross section over the same $R\simeq 3$–12 kpc range, indicating a segmented, non-coherent spiral morphology rather than a single large-scale grand-design pattern (see also Fig. 9).
Overall, our model does not conform to a grand design spiral archetype. Instead, while a spiral pattern is always present, it is mostly transient, with the gas and stars often out of sync. This mirrors observations of external galaxies, where gaseous spirals display greater variability due to the interstellar medium’s dynamical response (see e.g. Elmegreen and Elmegreen, 1987; Maschmann et al., 2024; Dobbs and Bonnell, 2008; Renaud et al., 2013). Such hybrid behaviour highlights the diversity of spiral structures under evolving galactic conditions (Foyle et al., 2011; Hart et al., 2016).
As noted in Sect. 1, studies suggest the Milky Way shows both grand design and multi-armed features. For example, Drimmel and Spergel (2001) and Churchwell et al. (2009) found a two-arm structure in the inner Galaxy, while the outer disc is more complex. Spitzer GLIMPSE observations confirm this duality, showing dominant and fainter arms (Benjamin et al., 2005). Gas surveys (Nakanishi and Sofue, 2003; Kalberla and Kerp, 2009) also reveal significant variation, consistent with our model. This suggests that the Milky Way’s spiral structure may be inherently transient, with gas responding in a complex, non-density-wave-like manner, a point we explore further in the following sections (4.2 and 5.1).
@@ 4.2 Solar neighbourhood
The recent work by Khanna et al. (2023) has identified significant streaming motions for stars within $3.5$ kpc from the Sun, demonstrating tentative correlations between the velocity structures and the spiral arms. Using an axisymmetric model to subtract from the observed velocities, they quantified overall velocity trends and found that gradients in the tangential and radial components ($\Delta(-V_{\phi}$) and $\Delta V_{R}$) were not uniform, with variations averaging $\sim 5$ km s-1 and $\sim 10$ km s-1 respectively. However, the extent to which such velocity patterns can be related to the dynamics of the stellar and gaseous spiral pattern in the MW is incredibly hard to ascertain from observations. Indeed, as mentioned in Section 1, our knowledge of the position and extent of the stellar and gaseous arms remains highly uncertain, making these results tentative at best.
@@
In order to shed some light into how the observed stellar kinematics pattern from Gaia may inform us of the underlying spiral arm structure and kinematics in the solar neighbourhood, here we investigate the equivalent stellar velocity patterns in our model, and then compare that to the underlying stellar and gaseous spiral pattern. For that, we obtain the velocity residuals in a similar way to the observations by Khanna et al. (2023) by contrasting the actual velocity fields observed in our simulation against an idealised rotation curve (see Appendix D for more details).
@@
In Figure 7, we present the heliocentric velocity residuals, $\Delta(-V_{\phi})$ and $\Delta V_{R}$, for the data from the Gaia DR3 study by Khanna et al. (2023) (top), alongside our model’s stellar (middle) and gaseous (bottom) components within the mid-plane (—zz— ¡ 0.25 kpc). Note that the stellar and gaseous panels use different colour-bar ranges to reflect their different dynamic ranges. On the Gaia data maps we include the spiral arm tracks from Reid et al. (2019), and in the simulation panels, we have included the stellar and gaseous spiral arm ridges (in black and grey respectively), as identified using an unsharpening technique (see full details in Appendix C). The gaseous ridges in the simulation broadly follow the Reid et al. (2019) tracks, and the general trends for the stellar kinematics in our model seem to be similar to those observed by Khanna et al. (2023), are qualitatively similar to the Gaia maps. The Sun lies near a transition between inward and outward motions in $\Delta V_{R}$ and in a region of moderately slower tangential velocities. The residual amplitudes are slightly larger in the model, reaching $\sim 15$ km s-1 in $\Delta(-V_{\phi})$ and $\sim 25$ km s-1 in $\Delta V_{R}$. Overall, the stellar kinematics in our simulation show a relatively weak correlation with the positions of the spiral ridges:neither $\Delta(-V_{\phi})$ nor $\Delta V_{R}$ show a systematic alignment with the stellar (black) or gaseous (grey) ridges across the map.
@@
Figure 7: Heliocentric velocity residual maps for $\Delta(-V_{\phi})$ (left) and $\Delta V_{R}$ (right) for Gaia DR3 data from Khanna et al. (2023) (top), and our model’s stellar (middle) and gaseous (bottom) components in the galactic mid-plane (—zz—¡ 0.25 kpc). The stellar maps use a binning identical to those of the Gaia data, of $250\times 250$ pc2 pixels, for a direct comparison, while the gas maps have been binned to cells of $60\times 60$ pc2. The solid grey lines on the top panels represent the gaseous spiral arm tracks from Reid et al. (2019). For our simulation, the stellar and gaseous spiral arm ridges extracted from the sharpened images of Fig. 15, are superimposed in black and grey respectively. The position of the sun is marked as a black circle, and the galactic centre is outside this image, at (8.275,0) kpc.
@@
Overall, our analysis shows that the pattern in the stellar structure and kinematics of our model at the chosen time is generally mild and consistent with Gaia observations of RGB stars, but that it translates into a gaseous pattern that is much sharper. This is also consistent with observations from Gaia Collaboration et al. (2023) who show that OB stars (which by being younger, are potentially better tracers of the underlying gas distribution that originated them) do align better with the gaseous spiral structures, specially towards the inner disc, but with low number densities their maps remain uncertain. Given the consistency between our results and MW observations, our model could therefore offer some insight into how the formation (and evolution) of the gaseous spiral structures of the MW, albeit potentially rapidly evolving, might be intricately inter-connected with the observed kinematics of the gas, and thus we explore this correlation further in the next section.